Andrew is 87 years younger than Katie. 9 years ago, Katie's age was 4 times Andrew's age. How old is Andrew now?

Found 2 solutions by stanbon, macston:
Answer by stanbon(75887)

Andrew is 87 years younger than Katie. 9 years ago, Katie's age was 4 times Andrew's age. How old is Andrew now?
----------
a = k-87
k-9 = 4(a-9)
-----
Substitute for "k" and solve for "a"::
(a+87-9) = 4a-36
-------
78 = 3a - 36
------
3a = 114
a = 38 (Andrew's age now)
---------------
Cheers,
Stan H.
------------

Answer by macston(5194)

.
A=Andrew's age; K=Kati's age=A+87yrs
.
K-9=4(A-9)
(A+87)-9=4A-36
A+78=4A-36
A+114=4A
114=3A
38=A
.
ANSWER; Andrew is 38 years old.
.
CHECK:
K=A+87=38+87=125
.
K-9=4(A-9)
125-9=4(38-9)
116=4(29)
116=116
.